AnakKrakatau1883:

Berniece solved the following equation: 4x − 2x + 8 = 6(x + 4) Step Work Justification 1 2x + 8 = 6(x + 4) Combine like terms 2 2x + 8 = 6x + 24 Distributive Property 3 −4x + 8 = 24 Subtraction Property of Equality 4 −4x = 16 Addition Property of Equality 5 x = −4 Division Property of Equality Which step has an incorrect justification? Step 1 Step 2 Step 3 Step 4

Vocaloid:

going from -4x + 8 = 24 to -4x = 16, they *subtracted* 8 from both sides. which property of equality is this?

AnakKrakatau1883:

Subtraction Property of Equality.

Vocaloid:

good so looking back at the original steps 3 and 4 3 −4x + 8 = 24 Subtraction Property of Equality 4 −4x = 16 Addition Property of Equality they actually used the subtraction property of equality to go from step 3 to 4, so step 4 should actually be subtraction property of equality, making step 4 w/ the incorrect justification, and thus the solution to the problem.

AnakKrakatau1883:

Thanks for helping!